1. What are the electric field and electric flux of a sphere with a radius of 0.500 m and carries a charGe of +1.00 uc at its center?

2. A long straight wire has a charge per unit length of 3.00 x 10^-12 C/m. At what distance from the wire is the electric field equal to 0.600 N/c.

3. How many excess electrons must be added to an isolated spherical conductor 0.100 meters in diameter to produce an electric field of 1800 N/C just outside the surface?

4. A closed irregular surface encloses a net charge of 1.50 x 10 C. What is the net electric flux through the surface?

5. What must be tge charge of an isolated spherical conductor 30.0 cm in diameter to produce an electric field of 150 N/C just outside the surface?


1
Expert's answer
2021-04-26T17:12:50-0400

1)


"\Phi=\frac{10^{-6}}{8.85\cdot10^{-12}}=1.13\cdot10^{5}\ Vm"

"E=\frac{1.13\cdot10^{5}}{4\pi(0.5)^2}=3.6\cdot10^{4}\frac{V}{m}"

2)

"\frac{3\cdot10^{-12}}{2\pi (8.85\cdot10^{-12})r}=0.6\\r=0.0899\ m"

3)


"n\frac{1.6\cdot10^{-19}}{4\pi(0.1)^2(8.85\cdot10^{-12})}=1800\\n=1.25\cdot10^{10}"

4)


"\Phi=\frac{15}{8.85\cdot10^{-12}}=1.69\cdot10^{12}\ Vm"

5)


"\frac{q}{4\pi(0.15)^2(8.85\cdot10^{-12})}=150\\q=3.75\cdot10^{-10}\ C"